OpenStudy (anonymous):

6. The cones below are similar, although not drawn to scale. Question Group 16 Question 2 What is the length of the radius of the larger cone? (1 point) 5 ft 6 ft 7 ft 8 ft

OpenStudy (anonymous):

\[\frac{ 15 }{ 5 }=\frac{ 18 }{ x }\]

OpenStudy (anonymous):

15x=90

OpenStudy (anonymous):

15x/15=x 90/15=6 x=6

OpenStudy (anonymous):

So, B is the answer.

OpenStudy (anonymous):

thank you